OpenStudy (anonymous):

-7 1/3 written as a fraction is?

OpenStudy (anonymous):

In these mixed numbers the first part, so in this case -7 is the whole number part than the next fraction 1/3 is added to the number. -7=-21/3 now add 1/3 -21/3 + 1/3=-20/3

OpenStudy (anonymous):

I never understand as a maths uni student that why this mixed number is important. It is not convenient to use it.